# Assembly Theory Reduced to Shannon entropy and Rendered Redundant by Naive Statistical Algorithms

### v7

@@ Acknowledgements
One regular argument made by the authors of Assembly Theory to defend their Assembly index against suggestions of similarities or full equivalence to algorithmic complexity is that their measure is computable, unlike algorithmic (Kolmogorov) complexity [81, 23, 1]. They have even provided a several-page mathematical proof to this effect. Clearly, the index did not need any proof because it is trivially computable, and this is now even more evident given their fully proven mathematical equivalence to popular lossless compression $\mathbf{LZ}$ algorithms that are upper bounded by Shannon entropy, which is trivially computable as well.
On the one hand, algorithmic (Kolmogorov) complexity is lower semi-computable, meaning it can be approximated from below, and some compression is sufficient proof of non-randomness. On the other hand, algorithmic complexity is semicomputable if the only answer accepted is the smallest computer program, but neither applications of algorithmic complexity nor the Assembly index are exclusively concerned with the shortest possible algorithm (or pathway). This means that finding any shorter pathway or computer program is computable. Moreover, for decades, researchers have introduced resource-bounded complexity measures based on or inspired by the same arguments of conciseness of mechanistic description, including our own Coding Theorem and Block Decomposition Methods [8, 7], which are fully computable, but also LZ(W) algorithms used in all sorts of applications for characterisation and identification of chemical and physical properties of life, selection and evolution [31, 30].
@@
In contrast, the measure we introduced over a decade ago [8, 7] combines the current state-of-the-art algorithms from which it draws its power or is based upon, does take into consideration all the possible pathways along which an object could have been constructed causally and mechanically, and gives proper weight to the many other pathways (not only the shortest one as in AT), thereby even implementing what the AT authors often claim to have implemented but did not. What they have done is aim directly for a statistical approximation of algorithmic complexity, with the known limitations of such approximations, which they present as features, limitations that all statistical Shannon-entropic indexes possess. All the while they’ve claimed to be different, even when such a claim is mathematically disproven  [23]. Indeed, they have not yet proven that they can do anything that Shannon entropy, Minimum Description Length (MDL) approaches or LZW and cognates can already do and have done for the last 50 years or more (e.g. [84]).
@@
This often arises from an academic culture that tends to prioritise the safe, well-defined boundaries of redundant methods for trivial problems over stronger methodological foundations. If we are to have any chance of dealing with causality, for example, there is no option but to deal with the uncomputability that arises from asking non-trivial questions about simple systems that are capable of universal computation [75, 9]. The fear of uncomputability reflects a narrowed scientific understanding of open-ended methods equipped to deal with open-ended challenges, such as the causal challenge of inverse problems.

### v10

@@ 6 Funding
One regular argument made by the authors of Assembly Theory to defend their Assembly index against suggestions of similarities or full equivalence to algorithmic complexity is that their measure is computable, unlike algorithmic (Kolmogorov) complexity [79, 26, 4]. They have even provided a several-page mathematical proof to this effect. Clearly, the index did not need any proof because it is trivially computable, and this is now even more evident given their fully proven mathematical equivalence to popular lossless compression LZ algorithms that are upper bounded by Shannon entropy, which is trivially computable as well.
On the one hand, algorithmic (Kolmogorov) complexity is lower semi-computable, meaning it can be approximated from below, and some compression is sufficient proof of nonrandomness. On the other hand, algorithmic complexity is semicomputable if the only answer accepted is the smallest computer program, but neither applications of algorithmic complexity nor the Assembly index are exclusively concerned with the shortest possible algorithm (or pathway). This means that finding any shorter pathway or computer program is computable. Moreover, for decades, researchers have introduced resource-bounded complexity measures based on or inspired by the same arguments of conciseness of mechanistic description, including our own Coding Theorem and Block Decomposition Methods [7, 6], which are fully computable, but also LZ(W) algorithms used in all sorts of applications for characterisation and identification of chemical and physical properties of life, selection and evolution [30, 11].
@@
In contrast, the measure we introduced over a decade ago [7, 6] combines the current state-of-the-art algorithms from which it draws its power or is based upon, does take into consideration all the possible pathways along which an object could have been constructed causally and mechanically, and gives proper weight to the many other pathways (not only the shortest one as in AT), thereby even implementing what the AT authors often claim to have implemented but did not. What they have done is aim directly for a statistical approximation of algorithmic complexity, with the known limitations of such approximations, which they present as features, limitations that all statistical Shannonentropic indexes possess. All the while they’ve claimed to be different, even when such a claim is mathematically disproven  [26]. Indeed, they have not yet proven that they can do anything that Shannon entropy, Minimum Description Length (MDL) approaches or LZW and cognates can already do and have done for the last 50 years or more (e.g. [82]).
@@
This often arises from an academic culture that tends to prioritise the safe, well-defined boundaries of redundant methods for trivial problems over stronger methodological foundations. If we are to have any chance of dealing with causality, for example, there is no option but to deal with the uncomputability that arises from asking nontrivial questions about simple systems that are capable of universal computation [72, 8]. The fear of uncomputability reflects a narrowed scientific understanding of open-ended methods equipped to deal with open-ended challenges, such as the causal challenge of inverse problems.
@@
Clustering Molecules from Experimental Data
Fig. S3 on the right-hand side, shows four different clustering algorithms to visualise how well the mass spectra peak matrices (i.e., m/z ratio vs. a number of peaks) cluster by the three categories. As shown, hierarchical clustering, Gaussian Mixture Models (GMM), KMeans cluster, SVM, Random Forest (RF), KNN, and a neural network (FFNN) or multilayer perceptron, all but PCA separate the three categories.
Figure S3: Clustering and machine learning classification performance on mass spectrometry peak matrices from source data showing that any naive statistical or machine learning algorithm (for both transparent- and opaque-box approaches) can reproduce the results that Assembly Theory deems the ones unifying biology and physics [2].
In Fig. S3, from A to D on the left-hand side, we see the ROC curves of four ML algorithms classifying the organic and inorganic molecules (N=30) from the Marshall et al. [1] paper’s Fig. 4 mass spectrometry mixtures into 3 categories: Dead, Inorganic, and Biological. SVM had the highest area under the curve (AUC), with 0.75, and a classification accuracy of 0.86. SVM suggests linear hyperplane separability.
The ML predictions are further supported by the ROC curves in Figure S4. Fig. S4 demonstrates that 1D-BDM and LZW compression measures significantly outperform MA as complexity measures in distinguishing between inorganic, dead, and biological molecules from [1] mass spectrometry data. Specifically, the AUC scores for BDM and LZW approaches are perfect or near-perfect (AUC = 1.00), while MA struggles with considerably lower AUC scores: within the 0.43-0.77 range with ML classification. This underperformance suggests that MA is suboptimal for detecting and quantifying biosignatures, possibly due to its less robust representation of molecular complexity in comparison to complexity measures from the AIT framework, like BDM and LZW, which more effectively capture the intricate patterns needed to differentiate between these categories. Consequently, MA (Ai) may not be well-suited for the study’s objective of identifying unique biosignatures of life based on molecular complexity and its multi-scale organisation.
Figure S4: ROC curves comparing mass spectrometry data of organic and inorganic molecules by way of comparing the measures of 1D-BDM and MA. LZW, and other compression schemes like RLE and Huffman, had similar values close to the evaluated ML performance of BDM (AUC scores of 0.89 for MLP and 1.00 for KNN and RF). It is clearly shown that BDM (and, similarly, other AIT measures) outperforms MA (the values attributed to the environmental samples’ mass spectrometry data in Marshall et al. [1] Figure 4) using basic machine learning algorithms. Likewise, LZW compression also outperforms MA, with a perfect AUC (1.00) using FFNN and RF and a better AUC than MA for KNN (0.69). MA has poor AUC scores of 0.47 for MLP and 0.43 in KNN and RF-based classification. All this even gives AT and MA the advantage of using their own small choice of molecules [1] rather than the exhaustive experiment we reported years before [22] on more than 15 000 compounds.
Therefore, these findings demonstrate that even naive ML approaches could achieve equal or superior classification performance compared to the assembly index (Ai), with SVM and FFNN achieving high precision and AUC scores, thus challenging the uniqueness and utility of Ai as a biosignature discriminator.